OpenStudy (halfdeafdarling):

okay for a question i had a minute ago I have created a function which looked like this (ill write it in the comments so I can use the symbols

OpenStudy (halfdeafdarling):

\[f(x)=1(x-1)^2+4\] This is a function we had to made using our birth date and being that I am january 4th it is fairly simple. With me doing that I went then to this: \[f(x)=1(x-1)(x-1)+4\]

OpenStudy (halfdeafdarling):

once I did this i distributed and got: \[f(x)=x^2-2x+5 \] would this be putting my function into standard form from it being in vertex form?

OpenStudy (halfdeafdarling):

@rebeccaxhawaii

rebeccaxhawaii (rebeccaxhawaii):

ax+bx=c this is standard form

OpenStudy (halfdeafdarling):

okay so from vertex form which i have above how would i go to standard?

OpenStudy (halfdeafdarling):

@Michele_Laino mind helping if you have a second?